OpenStudy (anonymous):

The lengths of the two legs of a right triangle are thirteen centimeters and ten centimeters. What is the exact length of the hypotenuse? A 3 centimeters B square root of ten centimeters C square root of sixty-nine centimeters D square root of two hundred sixty-nine centimeters

OpenStudy (jack1):

pythagoras theorm: \[h ^{2} = a^2 + b^2\]

OpenStudy (jack1):

a = 13, b = 10 so what does h = ???

OpenStudy (jack1):

@stepennd7

OpenStudy (anonymous):

\[(13 )^2 + ( 10 )^2 = x^2 \\160 + 100= x^2\\ x= \sqrt{269}\\x=16.40\]